OpenStudy (anonymous):

solve for x y=((2/5)x^2)+((14/10)x)+x^(2/5) solve for x y=((2/5)x^2)+((14/10)x)+x^(2/5) @Mathematics

OpenStudy (anonymous):

so you set it equal to 0 then solve for x

OpenStudy (turingtest):

are you asking or telling that?

OpenStudy (anonymous):

im asking

myininaya (myininaya):

are you trying to find x-intercepts

OpenStudy (turingtest):

not necessarily then, depends on what you need to know.

OpenStudy (anonymous):

no this is the derivative of a function im trying to find max and min

OpenStudy (turingtest):

take it away myininaya...

myininaya (myininaya):

so this is the derivative of some function do you want to write down the original function so we can check your derivative here?

myininaya (myininaya):

or you know its right?

OpenStudy (anonymous):

y = x^(2/5)((7/2)+x)

OpenStudy (anonymous):

just incase there is a easier way that I missed

myininaya (myininaya):

\[y'=\frac{2}{5}x^{\frac{2}{5}-1}(\frac{7}{2}+x)+x^\frac{2}{5}(0+1)\]

myininaya (myininaya):

\[y'=\frac{2}{5}x^{\frac{-4}{5}}(\frac{7}{2}+x)+x^\frac{2}{5}\]

OpenStudy (anonymous):

so that simplifies to (7(x+1))/5x^3/5 and then x=-1?

OpenStudy (turingtest):

that should be x^(-3/5) where it is x^(-4/5) above

OpenStudy (anonymous):

so X=-1 X=0

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

or is it not x=0

myininaya (myininaya):

\[y'=\frac{7}{5x^\frac{3}{5}}+\frac{2x}{5x^{\frac{3}{5}}}+x^\frac{2}{5}\] right since turingtest knows how to subtract and i don't my bad

OpenStudy (turingtest):

let myininaya do it, I'm eating and catching silly typos I know you can subtract

myininaya (myininaya):

\[y'=\frac{7+2x+5x}{x^\frac{3}{5}}\]

myininaya (myininaya):

please though let me know if i make an error

myininaya (myininaya):

\[y'=\frac{7(1+x)}{x^\frac{3}{5}}\]

myininaya (myininaya):

critical numbers or x=-1, 0

OpenStudy (turingtest):

looks good to me.

myininaya (myininaya):

great i need to go to sleep now

OpenStudy (turingtest):

g'night!

myininaya (myininaya):

night

OpenStudy (anonymous):

but 0 doesnt work

myininaya (myininaya):

0 is in the domain of the the function f f(0)=0